# Lower Bound of Nodal Sets in Elliptic Homogenization and Functions with Strong Maximum Principle

### v1

@@ 4 Special Case in SMP function
In this section, we will try to derive the lower bound in a different way and all the functions is defined in $\mathbb{R}^{2}$. It only needs the basic definition of Hausdorff measure. Also, it only requires the function to be continuous and has strong maximum principle. We also hope this argument will works for higher dimension. For convenience, we will use the following notation.
@@
We only need to prove if $u\in SMP\left(B_{1}\right),u\left(0\right)=0$, then
@@
where $T_{s,1}=\left\{s\leq|x|\leq 1\right\}$. WLOG, we assume that $\mathcal{H}^{1}(\left\{u=0\right\}\cap\overline{B_{1}})<\infty$. Otherwise, there is nothing to prove. We will prove the final results with the following two claims.
@@
For the second case, similarily, we have
@@
We rotate $\{x_{i}^{1}\}$ along the circle $\left\{r=|x_{i}^{1}|\right\}$ and also rotate $\{x_{i}^{2}\}$ along the circle $\left\{r=|x_{i}^{2}|\right\}$ to the points
@@
Then by the property of continuous function, we that know for any $s-\small{\frac{r_{s}}{2}\leq t<s}$
@@
$\mathcal{H}^{1}(\left\{u=0\right\}\cap\left\{s-\small{\frac{r_{s}}{2}}\leqslant r<s\right\})\geq\mathcal{H}^{1}(l_{1}\cup l_{2})=r_{s}$ (4.21)
As a result, we get
$\mathcal{H}^{1}(\left\{u=0\right\}\cap\left\{s-\small{\frac{r_{s}}{2}}\leqslant r\leqslant 1\right\})\geq 2(1-(s-\frac{r_{s}}{2}))$ (4.22)
@@
In fact we can pose weaker condition to $u$ as below
@@
By similar argument, we can derive the following theorem.
@@
We have the following equality holds

### v2

@@ 4 Special Case in SMP function
In this section, we will derive the lower bound in a different way and all the functions are defined in $\mathbb{R}^{2}$. It only needs the basic definition of Hausdorff measure. Also, it only requires the function to be continuous and has strong maximum principle. We also hope this argument will work for higher dimension. For convenience, we will use the following notation:
@@
We only need to prove that if $u\in SMP\left(B_{1}\right),u\left(0\right)=0$, then
@@
where $T_{s,1}=\left\{s\leq|x|\leq 1\right\}$. WLOG, we assume that $\mathcal{H}^{1}(\left\{u=0\right\}\cap\overline{B_{1}})<\infty$. Otherwise, there is nothing to prove. We will prove the final result with the following two claims.
@@
For the second case, similarly, we have
@@
We can rotate $\{x_{i}^{1}\}$ along the circle $\left\{r=|x_{i}^{1}|\right\}$ and also rotate $\{x_{i}^{2}\}$ along the circle $\left\{r=|x_{i}^{2}|\right\}$ to the points
@@
Then by the property of continuous function, we know that for any $s-\small{\frac{r_{s}}{2}\leq t<s}$
@@
$\mathcal{H}^{1}(\left\{u=0\right\}\cap\left\{s-\small{\frac{r_{s}}{2}}\leq r<s\right\})\geq\mathcal{H}^{1}(l_{1}\cup l_{2})=r_{s}$ (4.21)
Therefore, we obtain
$\mathcal{H}^{1}(\left\{u=0\right\}\cap\left\{s-\small{\frac{r_{s}}{2}}\leq r\leq 1\right\})\geq 2(1-(s-\frac{r_{s}}{2}))$ (4.22)
This finishes the proof of Claim 2. ∎
@@
In fact we can pose weaker condition to $u$ as below:
@@
By a similar argument, we can derive the following theorem:
@@
We have the following equality holds: